cos x + sin x = 1/2
Find the value of x.

Grade:12

1 Answers

bharat bajaj IIT Delhi
10 years ago
cos x + sin x = 1/2
multiply by 1/v2 on both sides :
1/v2 cos x + 1/v2 sin x = 1/2v2
cos x . cos 45 + sin x. sin 45 = 1/2v2
cos(x-45) = 1/2v2
x- 45 = cos^(-1) 1/2v2
Hence, x = 45 +cos^(-1) 1/2v2
Thanks
Bharat Bajaj
IIT Delhi
askiitians faculty